OpenStudy (anonymous):

the tongue of a chamelon is one and a half times its body length write this as a decimal

OpenStudy (anonymous):

1.5

sammixboo (sammixboo):

Well we have 1, and a 1/2. What is 1/2 in decimal form?

OpenStudy (anonymous):

1/2 = 0.5  = 3/2 = 1.5